OpenStudy (anonymous):

how do i solve Sin(2x)^2-Cos(2x)=Cos(2x)?

hartnn (hartnn):

how would you simplify and equation ? equation is solved expression is simplified maybe you need to solve ? try this... (sin 2x)^2 = 2 cos 2x

OpenStudy (anonymous):

@hartnn i meant that! I wrote simplify because thats how i read the problem in my head :) I'm trying to use double able identities to solve it..

OpenStudy (anonymous):

do i use the identity Sin2x=2Sinx*cosx to solve for the left side?

hartnn (hartnn):

so, solving.... yes, use sin 2x = 2 sin x cos x 4 sin^2 x cos^2x = 2( cos^2x - sin^2 x) now replace all sin^2 x by (1-cos^2x) [or all cos^2x = 1-sin^2 x] and get everything in terms of cos^2 x, you will get a quadratic in cos^2 x which you can easily solve